When 2 times a number is subtracted from 5, the absolute value of the difference is at least 6. Use interval notation to express the set of all numbers that satisfy this condition

Found 3 solutions by mananth, ikleyn, greenestamps:
Answer by mananth(16946)

When 2 times a number is subtracted from 5, the absolute value of the difference is at least 6. Use interval notation to express the set of all numbers that satisfy this condition

| 5-2x | <=6

5-2x <= -6 or 5-2x >= 6
5-2x-5 <= -6 -5 or 5-2x -5 >= 6-5
-2x <= -11 or -2x >= 1
-2x/-2 >= -11/-2 or -2x/-2 <= 1/-2
x >= 11/2 or x <= -1/2
(-infinity, -1/2]) u [11/2, infinity)

Answer by ikleyn(52776)
Answer by greenestamps(13198)



(Note for the other tutors who have responded....)

The statement of the problem is that the absolute value of the difference is AT LEAST 6 -- not at most 6.